OpenStudy (anonymous):

Please help?? 28g-T=28a T-15g=15a

OpenStudy (anonymous):

What do you need help with?.. can't u solve the linear equations? .. u need help with solving them??

OpenStudy (anonymous):

resolve this issue

OpenStudy (anonymous):

just add the two equations

OpenStudy (anonymous):

can explain step by step?

OpenStudy (anonymous):

yea.. the first step is adding the two equations.. add and tell me what happens on left side and what happens on right side?

OpenStudy (anonymous):

that's the problem, I could do until that point. now do not know what to do anymore ..

OpenStudy (anonymous):

can't u do addition man? :P

OpenStudy (anonymous):

28g-T=28a (1) T-15g=15a (2) (1)+(2)=13g=43a a=13g/43